Bear3:

The equation y=19(1.057)x represents book sales in billions of dollars since 1990. Let x represent the number of years. Let x = 0 represent book sales in 1990. What were the total sales, in billions of dollars, in 2010?

AZ:

\( y = 19 (1.057)^x \) x = 0 is the book sales in 1990 and we want to calculate the book sales in 2010 so what is 2010 - 1990 = ? that will be the x value we use to put into the formula to solve for y, the total sale in billions of dollars so what is 2010 - 1990 =

AZ:

Once you find that number, just take a calculator and do 1.057^(that number) and then multiply by 19 and that will give you the total sales in 2010